OpenStudy (anonymous):

A pair of equations is shown below. x + y = 2 y = one halfx + 5 If the two equations are graphed, at what point do the lines representing the two equations intersect? (4, −2) (−2, 4) (2, 5) (5, −2) Please help I will fail if I dont do this plese help me

OpenStudy (butterflydreamer):

firstly, you want to make y the subject in the first equation " x + y = 2" :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Then what

OpenStudy (butterflydreamer):

so can you tell me what it would be?

OpenStudy (anonymous):

I know you isolate it but I do not really understand

OpenStudy (butterflydreamer):

\[x + y = 2 \rightarrow y = -x + 2 \] right? All i did was subtract "x" from both sides to make y the subject :)

OpenStudy (anonymous):

ok then what

OpenStudy (anonymous):

you do it to x also

OpenStudy (butterflydreamer):

so now we have two equations: